OpenStudy (anonymous):

Two dice are rolled. What is the probability that the sum of the dice is bigger than 5 or that at least one of the dice shows an odd number?

OpenStudy (anonymous):

total possible outcomes r 36 sum bigger than 5 is possible in 26 outcomes So P(sum is bigger than 5) = 26/36 = 13/18

OpenStudy (anonymous):

Thanks, can u pls explain the workings to achieve these answers?

OpenStudy (anonymous):

P(at least one dice shows an odd number) = 27/36

OpenStudy (anonymous):

You can do this easily by writing out all the possible outcomes as follows: 11 12 13 14 15 16 21 22 23 24 25 26 31 32 33 34 35 36 41 42 43 44 45 46 51 52 53 54 55 56 61 62 63 64 65 66 Now u can check for yourself....

OpenStudy (anonymous):

pls click the GOOD ANSWER button if u find the answer satisfactory so that the question appears as solved and closed